OpenStudy (anonymous):

Solve by elimination 4x - y = 29 6x - 2y = 40

OpenStudy (anonymous):

Do i multiply by 2 to get the y's to be the same

OpenStudy (anonymous):

just multiply first equation bu -1

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

-2 sorry

OpenStudy (anonymous):

-8x-2y=-58

OpenStudy (anonymous):

2 by -24 is -48

OpenStudy (anonymous):

then add two equation to get rid of -2y

OpenStudy (anonymous):

??????? i'm so confused one person plz

OpenStudy (anonymous):

-8x=-58 6x=40

OpenStudy (anonymous):

-2x=-18 x=9

OpenStudy (anonymous):

????

OpenStudy (anonymous):

-2(4x - y = 29) 6x - 2y = 40 ----------------- -8x + 2y = -58 6x - 2y = 40 ----------------- -2x + 0y = -18 x = 9 now substitute x into any of the two equations to find y. 6(2) -2y = 40 12 - 2y = 40 -2y = 40 - 12 -2y = 28 y = -14

OpenStudy (anonymous):

thx for the help

OpenStudy (anonymous):

Sorry,made a mistake. 6(9) - 2y = 40 54 - 2y = 40 -2y = -14 y = 7

OpenStudy (anonymous):

Check to see if x = 9 and y =7 are solutions to your problem. Plug it into 4x - y = 29 4(9) - 7 = 29 36 - 7 = 29 29 = 29